OpenStudy (christos):

Give me the first step to solve (e^x - e^-x) / 2 = 2

OpenStudy (anonymous):

the first step is to multiply both sides by e^x

OpenStudy (anonymous):

then the next step make a proper substitution

OpenStudy (anonymous):

\[[e^x-e^{-x}=4]e^x\]

OpenStudy (anonymous):

\[e^{2x}-4e^x-1=0\]

OpenStudy (anonymous):

let m=e^x

OpenStudy (anonymous):

\[m^2-4m-1=0\]

OpenStudy (anonymous):

then find the values of m, hence take the natural log to find the value of x

OpenStudy (raden):

to get the value of m, use quadratic formula... it is cant be factored

OpenStudy (anonymous):

yes this basic of how to solve a quadratic equation must be known by students, otherwise they must not be studying maths....